OpenStudy (anonymous):

solve for x: √2x=3=x

OpenStudy (cwrw238):

thers must be a mistake in yur question jake - theres two '=' signs

OpenStudy (anonymous):

oh sorry its √2x+3 =x

OpenStudy (slaaibak):

\[\sqrt{2x} or \sqrt2 x\] ?

OpenStudy (slaaibak):

or\[\sqrt{2x+3}\] ?

OpenStudy (anonymous):

\[\sqrt{2x+3}=x\]

OpenStudy (slaaibak):

Cool. Square both sides to get: 2x + 3 = x^2 x^2 - 2x -3 = 0 (x+1)(x-3) = 0 x = -1 or x=3

OpenStudy (cwrw238):

square both sides

OpenStudy (slaaibak):

Now, test the answers. It can't be -1 So answer is 3.

OpenStudy (slaaibak):

^No. It can't be -1 because: sqrt(1) is not equal to -1. Always plug in the values and test it.